Changed Passwords for Schema Owners and modified related database links from As1012p, Crm11ip, Banner. Aslo, modified BJP’s parameter file.
CRM was temporarily unavailable for for about 40 minutes.
At approximately 10:30 AM, Internet service was resumed as Integra Telecom moved our routes to a temporary router. Internet service resumed until about 12:30 PM, when Integra moved our routes back to our normal router, causing a brief 5-10-minute interruption in service.
Updated Millennium DB Parameters to improve I/O performance.
SGA_MAX_SIZE, SHARED_POOL_SIZE, DB_CACHE_SIZE
A major fiber feed for Integra has been broken. A pole damaged along route 167 has broken the fiber feed for the state of Washington. This feed cannot be repaired until later this afternoon when 167 can be shutdown to restring the fiber across the highway. There is no estimated time of repair.
Updated Cascade DB global parameters to help performance:
Updated Millennium DB global parameters to help performance:
Updated CRM database global parameter to allow for better performance monitoring.
Email sent to liaisons:
Thursday April 17, Information Services will be performing critical changes on our database systems to improve recent performance concerns.
Access to all database systems including Cascade, Cascade Web, Banner, Famis, Basis and Millennium will be unavailable from 5:30 AM until 6 AM Thursday.
If you have questions or concerns regarding this outage, please contact Jenifer Brauen by telephone at ext. 2725, 253.879.2725, or via email at firstname.lastname@example.org. Thank you for understanding our need to conduct critical systems work.
– Jenifer Brauen
We have deployed wireless throughout Jones Hall. It is currently active and available to users.
Brought Millennium application off-line, restarted the Oracle database instance and restarted the application. This was done by the recommendation of Sage in coordination with database connectivity changes on 4.11.08.
Our primary database servers providing support for
- Cascade Web
experienced high load April 12 – 14 resulting in performance concerns and periodic service interruptions. Regular service levels were restored by 5pm Monday, April 14th.
The high load conditions were caused by new data load and warehousing load activity.
Upgraded the Oracle Client tools to 9i and removed the 8i Client files.
Beginning April 1st, the Mailman listservs started failing to deliver messages to their membership. This happened because one of the real-time blackhole lists that Mailman was subscribed to went out of service permanently, causing Mailman’s mail server (sendmail) to reject all messages outgoing to the list memberships. So the messages were received by Mailman and archived properly, but they were not delivered. This was found and resolved on April 7th.
This coincided with another email problem on April 2-4, which masked the listserv issue, which is why it took so long to find and resolve.